# MLC at Feigenbaum points

### v2

@@ 2.3.2. Superattracting model
  • components of ${\mathfrak{T}}_{f}\setminus f^{-m}\left({\boldsymbol{\mathfrak{B}}}^{[n]}_{f}\right)$ and components of ${\mathfrak{T}}_{\circ}\setminus f^{-m}_{\circ}\left({\boldsymbol{\mathfrak{B}}}^{[n]}_{\circ}\right)$ for all $m\geq 0$; and
  • components of $f^{-m}\left({\boldsymbol{\mathfrak{B}}}^{[n]}_{f}\right)$ and components of $f^{-m}_{\circ}\left({\boldsymbol{\mathfrak{B}}}^{[n]}_{\circ}\right)$ for all $m\geq 0$
@@
${\mathfrak{B}}^{[1]}_{a}$${\mathfrak{B}}^{[1],(6p_{2})}_{a}$${\mathfrak{B}}^{\prime\prime}_{1}$$\ell_{1}$${\mathfrak{B}}^{\prime}_{1}$${\mathfrak{B}}_{i}^{[2]}$${\mathfrak{B}}_{i+1}^{[2]}$${\mathfrak{B}}^{\prime\prime}_{2}$${\mathfrak{B}}^{\prime}_{2}$$\ell_{2}$${\mathfrak{B}}^{[1]}_{b}$$\alpha$
@@ 2.3.3. Satellite combinatorics
Suppose that $f$ is at least 33 33 times renormalizable and the first renormalization of $f$ is satellite. Then all ${\mathfrak{K}}_{i}^{[1]}$ are organized in the satellite flower around the $\alpha$ fixed point of $f$:
[…]
@@
For a satellite $f$ as above, let ${\mathfrak{B}}^{[1]}_{a},{\mathfrak{B}}^{[1]}_{b}$ be its two level 11 11 bushes. Then the continuum ${\mathfrak{T}}^{(6p_{2})}_{f}$ contains geodesic continua $\ell_{1}$ and $\ell_{2}$ (as in §2.2.2) connecting preperiodic lifts ${\mathfrak{B}}^{\prime}_{1},{\mathfrak{B}}^{\prime\prime}_{1}$ and ${\mathfrak{B}}^{\prime}_{2},{\mathfrak{B}}^{\prime\prime}_{2}$ of ${\mathfrak{B}}^{[1]}_{b}$ of preperiods $\leq 6p_{2}$ such that, see Figure 6:
@@
  • the geodesic continuum $\ell\subset{\mathfrak{T}}_{f}$ connecting $\ell_{1},\ell_{2}$ intersects little preperidic bushes ${\mathfrak{B}}_{i}^{[2]},{\mathfrak{B}}^{[2]}_{i+1}$of level 22 22 with preperiods less than $6p_{1}$;
  • ${\mathfrak{B}}_{i}^{[2]},{\mathfrak{B}}^{[2]}_{i+1}$ are disjoint from $\ell_{1}\cup\ell_{2}$.
We remark that $\ell,{\mathfrak{B}}_{i}^{[2]},{\mathfrak{B}}^{[2]}_{i+1}$ are within ${\mathfrak{B}}^{[1],(6p_{2})}_{a}$.
@@
Let ${\mathfrak{B}}_{\gamma}^{[2]}\subset{\mathfrak{B}}^{[1]}_{a}$ be the level 22 22 periodic bush closest to $\alpha$; i.e. $({\mathfrak{T}}_{f}\cap{\mathfrak{B}}^{[1]}_{a})\setminus{\mathfrak{B}}^{[2]}_{\gamma}$ and $\alpha$ are in different components of ${\mathfrak{T}}_{f}\setminus{\mathfrak{B}}_{\gamma}^{[2]}$. Let the ${\mathfrak{B}}_{t}^{[2]}\subset{\mathfrak{B}}^{[1],(tp_{1})}_{a}$ be the lifts of ${\mathfrak{B}}_{\gamma}^{[2]}$ under $f^{tp_{1}}$ towards $\alpha$; i.e. each ${\mathfrak{B}}_{t}^{[2]}$ separate $\alpha$ from ${\mathfrak{B}}^{[1],(tp_{1}-p_{1})}_{a}$.
Since $f^{6p_{2}}$ has critical points in ${\mathfrak{B}}^{[2]}_{1}$ and in ${\mathfrak{B}}_{4}^{[2]}$, we can select $\ell_{1}$ and $\ell_{2}$ passing through these critical points such that $\ell_{1},\ell_{2}$ connect preperiodic lifts ${\mathfrak{B}}^{\prime}_{1},{\mathfrak{B}}^{\prime\prime}_{1}$ and ${\mathfrak{B}}^{\prime}_{2},{\mathfrak{B}}^{\prime\prime}_{2}$ of ${\mathfrak{B}}^{[1]}_{b}$. Then $\ell_{1},\ell_{2}$ separate ${\mathfrak{B}}^{[1]}_{a}$ from $\alpha$, and $\ell_{1}\cup\ell_{2}$ separate ${\mathfrak{B}}^{[2]}_{2},{\mathfrak{B}}^{[2]}_{3}$ from $\alpha$ and ${\mathfrak{B}}^{[1]}_{a}$; i.e. we can take ${\mathfrak{B}}^{[2]}_{i},{\mathfrak{B}}^{[2]}_{i+1}$ to be ${\mathfrak{B}}^{[2]}_{2},{\mathfrak{B}}^{[2]}_{3}$. ∎
@@ 2.4.2. Arc diagrams of ql maps
(In other words, the $\ell_{i}$ are contractible into ${\boldsymbol{\mathfrak{B}}}^{[n],(1)}$.) Figure 7 gives an example of an invariant arc diagram.
@@
Write $\alpha_{i}=[\gamma_{i}]$ and consider $Y^{\prime}\coloneqq Y\setminus\left({\boldsymbol{\mathfrak{B}}}^{[n]}\cup\Gamma\right)$, where $\Gamma=\bigcup_{i}\gamma_{i}$. Then one of the little bushes ${\mathfrak{B}}^{[n]}_{i}$ is accessible from the outermost component of $\partial Y^{\prime}$. Therefore, we can select a proper path $\ell\subset Y^{\prime}$ connecting $\partial Y$ and ${\boldsymbol{\mathfrak{B}}}^{n}$. For every such a path $\ell$, its legal pullback $\widetilde{\ell}$ is any of its lift connecting ${\boldsymbol{\mathfrak{B}}}^{[n]}$ and $\partial X$ and concatenated by a path in $Y\setminus X$ so that ~$\widetilde{\ell}\subset Y\setminus{\boldsymbol{\mathfrak{B}}}^{[n]}$ is a proper path connecting $\partial Y$ and ${\boldsymbol{\mathfrak{B}}}^{n}$. Since $A$ is invariant, $\widetilde{\ell}$ can be again represented in $Y^{\prime}$ up to a proper homotopy in $Y\setminus{\boldsymbol{\mathfrak{B}}}^{[n]}$.
@@ 2.4.3. Genuine periodic arcs
  • every component of $\ell_{i}\setminus{\boldsymbol{\mathfrak{B}}}^{[n],(m)}$ is trivial in $X\setminus{\boldsymbol{\mathfrak{B}}}^{[n],(m)}$ (with respect to a proper homotopy, see §2.4.1) .
@@ 2.4.4. Arcs inside and outside of satellite flowers
such that $\alpha$ overflows $\alpha_{1},\alpha_{2}$ in the following sense: any decomposition 2.11 of $\alpha=[\gamma]$ rel ${\boldsymbol{\mathfrak{B}}}^{[n],(2p_{n})}$ contains two paths $\gamma_{a},\gamma_{b}$ representing $\alpha_{1}=[\gamma_{a}]$ and $\alpha_{2}=[\gamma_{b}]$.
@@ 2.4.5. Inhomogeneous configuration
An AD $A=\{\alpha_{i}\}$ on $X\setminus{\boldsymbol{\Upsilon}}$ is horizontal if every $\alpha_{i}$ connects components of ${\boldsymbol{\Upsilon}}$. A horizontal AD $A=\{\alpha_{i}=[\gamma_{i}]\}$ is $f^{p_{1}}$-invariant if every $\alpha_{i}$ can be represented up to a proper homotopy in $X\setminus{\boldsymbol{\Upsilon}}$ in an arbitrary small neighborhood of $f^{-p_{1}}\left({\boldsymbol{\Upsilon}}\cup\Gamma\right)$, where $\Gamma=\bigcup_{i}\gamma_{i}$. By replacing $f$ with its superattracting model (§2.3.2, compare also wit Lemma 2.2), we obtain:
@@ 2.5. Teichmüller contraction
be the ql straightening map. Let us remove from every satellite ${\mathcal{M}}_{i}$ a small open neighborhood of its cusp such that the remaining ${\mathcal{M}}^{\circ}_{i}$ contains all the preimages of the ${\mathcal{M}}_{j}$ under $R\colon{\mathcal{M}}_{i}\to{\mathcal{M}}$. For a primitive copy ${\mathcal{M}}_{i}$, set ${\mathcal{M}}_{i}^{\circ}\coloneqq{\mathcal{M}}_{i}$. By compactness, there is a $\Delta>1$ such that every $f_{c}$ with $c\in\bigcup_{i}{\mathcal{M}}_{i}^{\circ}$ has a ql restriction $g_{1}=f_{c}^{p_{1}}\colon X_{1,c}\to Y_{1,c}$ around its maximal little Julia set containing 00 0 such that $g_{1}$ satisfies (II) and such that $g_{1}$ is hybrid conjugate via $h_{c}\colon Y_{1,1}\to Y_{R(c)}$ with dilatation $\leq\Delta$ to a ql restriction $f_{R(c)}\colon X_{R(c)}\to Y_{R(c)}$. Moreover, we can select the $Y_{1,c}$ and $Y_{c}$ so that $Y_{c}\Supset Y_{1,c}$.
@@
$\Upsilon\coloneqq\bigcup_{m\leq n}\left({\boldsymbol{\mathfrak{B}}}^{[m]}_{f}\setminus{\mathfrak{B}}^{[m]}_{f,0}\right).$
@@ 3. $\psi^{\bullet}$-ql renormalization and near-degenerate regime
Given a compact subset $K\Subset S$, we denote by ${\mathcal{F}}(S,K)$ the family of non-trivial proper curves in $S\setminus K$ emerging from $\partial S$. We write
@@ 3.1. Outline
The correspondence $F_{n,i}$ in (3.3) is called a $\psi^{\bullet}$-ql map. It consists of a degree 22 22 branched covering $f_{n,i}$ and an immersion $\iota_{n,i}$, see definitions in §3.4.2. We define the Julia set and bush of $F_{n,i}$ to be that of the ql map $f_{n,i}$ (3.2):
@@
A $\psi^{\bullet}$-ql renormalization can be naturally defined for a $\psi^{\bullet}$-ql map; i.e., $\psi^{\bullet}$-ql renormalization can be iterated. Sup-Chain Rule for renormalization domains is stated in §3.4.4.
@@
by considering the fiber product §3.4.6. We have a natural $1\displaystyle 2^{k}:$ 1 branched covering $f^{k}\colon U^{k}\to V$ and an immersion $\iota^{k}\colon U^{k}\to V$.
@@ A Euclidean rectangle is a rectangle $E_{x}\coloneqq[0,x]\times[0,1]\subset{\mathbb{C}}$, where:
Assume that ${\mathcal{W}}(E_{x})>2$. The left and right 11 11-buffers of $E_{x}$ are defined
@@
is obtained from $E_{x}$ by removing 11 11-buffers. If ${\mathcal{W}}(E_{x})\leq 2$, then we set $E_{x}^{\operatorname{new}}\coloneqq\emptyset$. Similarly, buffers of any width are defined.
@@ 3.2.1. Monotonicity and Grötzsch inequality
We say a family of curves $SS\displaystyle\SS$ overflows a family ${\mathcal{G}}$ if every curve $SS\displaystyle\gamma\in\SS$ contains a subcurve $\gamma^{\prime}\in{\mathcal{G}}$. We also say that
@@ 3.3.1. Proper rectangles
A proper rectangle $R$ in an open hyperbolic surface $S$ is a rectangle in its ideal completion $\overline{S}=S\cup\partial^{i}S$ such that $\partial^{h}R\subset\partial^{i}S$. We naturally view vertical curves of $R$ as proper (and open) paths in $S$. We say that $R$ connects boundary components $J_{0},J_{1}\subset\partial^{i}R$ if $\partial^{h,0}R\subset J_{0},\ \partial^{h,1}R\subset J_{1}$.
@@ 3.3.2. Universal cover
For compact Riemann surface $V$ as in §2.4.1x, write its non-empty boundary as $\partial V=\bigsqcup_{k}J_{k}$. We set $\pi:\mathbb{D}\rightarrow\operatorname{int}V$ to be the universal covering, and we let ${\Lambda}\subset{\mathbb{T}}=\partial\mathbb{D}$ be the limit set for the group ${\Delta}$ of deck transformations of $\pi$. Then $\pi$ extends continuously to ${\mathbb{T}}\setminus{\Lambda}$, and $\pi:\overline{\mathbb{D}}\setminus{\Lambda}\rightarrow V$ is the universal covering of $V$. Moreover, $\pi$ restricted to any component of ${\mathbb{T}}\setminus{\Lambda}$ gives us a universal covering of some component $J_{k}$ of $\partial V$. Such a component of ${\mathbb{T}}\setminus{\Lambda}$ will be denoted ${\mathbf{J}}_{k}$ (usually we will need only one component for each $k$, so we will not use an extra label for it). The stabilizer of ${\mathbf{J}}_{k}$ in ${\Delta}$ is a cyclic group generated by a hyperbolic Möbius transformation $\tau_{k}$ with fixed points in $\partial{\mathbf{J}}_{k}$.

### v3

@@ 2.3.2. Superattracting model
[…]
@@
${\mathfrak{B}}^{[1]}_{a}$${\mathfrak{B}}^{[1],(6p_{2})}_{a}$${\mathfrak{B}}^{\prime\prime}$$\ell_{1}$${\mathfrak{B}}^{\prime}$${\mathfrak{B}}_{i_{1}}^{[2]}$${\mathfrak{B}}_{i_{2}}^{[2]}$${\mathfrak{B}}^{\prime\prime\prime}$${\mathfrak{B}}^{\text{iv}}$$\ell_{2}$${\mathfrak{B}}^{[1]}_{b}$$\alpha$
@@ 2.3.3. Satellite combinatorics
Suppose that $f$ is at least $3$ times renormalizable and the first renormalization of $f$ is satellite. Then all ${\mathfrak{K}}_{i}^{[1]}$ are organized in the satellite flower around the $\alpha$ fixed point of $f$:
${\mathfrak{B}}_{f}={\boldsymbol{\mathfrak{K}}}^{[1]}_{f}\hskip 28.45274pt\text{ because }\hskip 28.45274pt{\boldsymbol{\mathfrak{K}}}^{[1]}_{f}\supset{\mathfrak{T}}_{f},$
@@
For a satellite $f$ as above, let ${\mathfrak{B}}^{[1]}_{a},{\mathfrak{B}}^{[1]}_{b}$ be two of its level $1$ periodic bushes. Then the continuum ${\mathfrak{T}}^{(6p_{2})}_{f}$ contains geodesic continua $\ell_{1}$ and $\ell_{2}$ (as in §2.2.2) connecting preperiodic lifts ${\mathfrak{B}}^{\prime},{\mathfrak{B}}^{\prime\prime}$ and ${\mathfrak{B}}^{\prime\prime\prime},{\mathfrak{B}}^{\text{iv}}$ of ${\mathfrak{B}}^{[1]}_{b}$ of preperiods $\leq 6p_{2}$ such that, see Figure 6:
@@
  • the geodesic continuum $\ell\subset{\mathfrak{T}}_{f}$ connecting $\ell_{1},\ell_{2}$ intersects little preperiodic bushes ${\mathfrak{B}}_{i_{1}}^{[2]},{\mathfrak{B}}^{[2]}_{i_{2}}$ of level $2$ with preperiods less than $6p_{1}$;
  • ${\mathfrak{B}}_{i_{1}}^{[2]},{\mathfrak{B}}^{[2]}_{i_{2}}$ are disjoint from $\ell_{1}\cup\ell_{2}$.
We remark that $\ell,{\mathfrak{B}}_{i_{1}}^{[2]},{\mathfrak{B}}^{[2]}_{i_{2}}$ are within ${\mathfrak{B}}^{[1],(6p_{2})}_{a}$.
@@
Let ${\mathfrak{B}}_{\gamma}^{[2]}\subset{\mathfrak{B}}^{[1]}_{a}$ be the level $2$ periodic bush closest to $\alpha$; i.e. $({\mathfrak{T}}_{f}\cap{\mathfrak{B}}^{[1]}_{a})\setminus{\mathfrak{B}}^{[2]}_{\gamma}$ and $\alpha$ are in different components of ${\mathfrak{T}}_{f}\setminus{\mathfrak{B}}_{\gamma}^{[2]}$. Let the ${\mathfrak{B}}_{i(t)}^{[2]}\subset{\mathfrak{B}}^{[1],(tp_{1})}_{a}$ be the lifts of ${\mathfrak{B}}_{\gamma}^{[2]}$ under $f^{tp_{1}}$ towards $\alpha$; i.e. each ${\mathfrak{B}}_{i(t)}^{[2]}$ separates $\alpha$ from ${\mathfrak{B}}^{[1],(tp_{1}-p_{1})}_{a}$.
Since $f^{6p_{2}}$ has critical points in ${\mathfrak{B}}^{[2]}_{i(1)}$ and in ${\mathfrak{B}}_{i(4)}^{[2]}$, we can select $\ell_{1}$ and $\ell_{2}$ passing through these critical points such that $\ell_{1},\ell_{2}$ connect preperiodic lifts ${\mathfrak{B}}^{\prime},{\mathfrak{B}}^{\prime\prime}$ and ${\mathfrak{B}}^{\prime\prime\prime},{\mathfrak{B}}^{\text{iv}}$ of ${\mathfrak{B}}^{[1]}_{b}$. Then $\ell_{1},\ell_{2}$ separate ${\mathfrak{B}}^{[1]}_{a}$ from $\alpha$, and $\ell_{1}\cup\ell_{2}$ separate ${\mathfrak{B}}^{[2]}_{i(2)},{\mathfrak{B}}^{[2]}_{i(3)}$ from $\alpha$ and ${\mathfrak{B}}^{[1]}_{a}$; i.e. we can take ${\mathfrak{B}}^{[2]}_{i_{1}},{\mathfrak{B}}^{[2]}_{i_{2}}$ to be ${\mathfrak{B}}^{[2]}_{i(2)},{\mathfrak{B}}^{[2]}_{i(3)}$. ∎
@@ 2.4.2. Arc diagrams of ql maps
(In other words, the $\ell_{i}$ are contractible into ${\boldsymbol{\mathfrak{B}}}^{[n],(1)}$.) Here $f^{*}(A)$ is the pullback of $A$. Figure 7 gives an example of an invariant arc diagram.
@@
Write $\alpha_{i}=[\gamma_{i}]$ and consider $Y^{\prime}\coloneqq Y\setminus\left({\boldsymbol{\mathfrak{B}}}^{[n]}\cup\Gamma\right)$, where $\Gamma=\bigcup_{i}\gamma_{i}$. Then one of the little bushes ${\mathfrak{B}}^{[n]}_{i}$ is accessible from the outermost component of $\partial Y^{\prime}$. Therefore, we can select a proper path $\ell\subset Y^{\prime}$ connecting $\partial Y$ and ${\boldsymbol{\mathfrak{B}}}^{[n]}$. For every such a path $\ell$, its legal pullback $\widetilde{\ell}$ is any of its lift connecting ${\boldsymbol{\mathfrak{B}}}^{[n]}$ and $\partial X$ and concatenated by a path in $Y\setminus X$ so that $\widetilde{\ell}\subset Y\setminus{\boldsymbol{\mathfrak{B}}}^{[n]}$ is a proper path connecting $\partial Y$ and ${\boldsymbol{\mathfrak{B}}}^{n}$. Since $A$ is invariant, $\widetilde{\ell}$ can be again represented in $Y^{\prime}$ up to a proper homotopy in $Y\setminus{\boldsymbol{\mathfrak{B}}}^{[n]}$.
@@ 2.4.3. Genuine periodic arcs
  • every component of $\ell_{i}\setminus{\boldsymbol{\mathfrak{B}}}^{[n],(m)}$ is trivial in $X^{m}\setminus{\boldsymbol{\mathfrak{B}}}^{[n],(m)}$ (with respect to a proper homotopy, see §2.4.1) .
@@ 2.4.4. Arcs inside and outside of satellite flowers
such that $\alpha$ overflows $\alpha_{1},\alpha_{2}$ in the following sense: any decomposition (2.11) of $\alpha=[\gamma]$ rel ${\boldsymbol{\mathfrak{B}}}^{[n],(2p_{n})}$ contains two paths $\gamma_{a},\gamma_{b}$ representing $\alpha_{1}=[\gamma_{a}]$ and $\alpha_{2}=[\gamma_{b}]$.
@@ 2.4.5. Inhomogeneous configuration
An AD $A=\{\alpha_{i}\}$ on $X\setminus{\boldsymbol{\Upsilon}}$ is horizontal if every $\alpha_{i}$ connects components of ${\boldsymbol{\Upsilon}}$. A horizontal AD $A=\{\alpha_{i}=[\gamma_{i}]\}$ is $f^{p_{1}}$-invariant if every $\alpha_{i}$ can be represented up to a proper homotopy in $X\setminus{\boldsymbol{\Upsilon}}$ in an arbitrary small neighborhood of $f^{-p_{1}}\left({\boldsymbol{\Upsilon}}\cup\Gamma\right)$, where $\Gamma=\bigcup_{i}\gamma_{i}$. By replacing $f$ with its superattracting model (§2.3.2, compare also with Lemma 2.2), we obtain:
@@ 2.5. Teichmüller contraction
be the ql straightening map. Let us remove from every satellite ${\mathcal{M}}_{i}$ a small open neighborhood of its cusp such that the remaining ${\mathcal{M}}^{\circ}_{i}$ contains all the preimages of the ${\mathcal{M}}_{j}$ under $R\colon{\mathcal{M}}_{i}\to{\mathcal{M}}$. For a primitive copy ${\mathcal{M}}_{i}$, set ${\mathcal{M}}_{i}^{\circ}\coloneqq{\mathcal{M}}_{i}$. By compactness, there is a $\Delta>1$ such that every $f_{c}$ with $c\in\bigcup_{i}{\mathcal{M}}_{i}^{\circ}$ has a ql restriction $g_{1}=f_{c}^{p_{1}}\colon X_{1,c}\to Y_{1,c}$ around its maximal little Julia set containing $0$ such that $g_{1}$ satisfies (II) and such that $g_{1}$ is hybrid conjugate via $h_{c}\colon Y_{1,1}\to Y_{R(c)}$ with dilatation $\leq\Delta$ to a ql restriction $f_{R(c)}\colon X_{R(c)}\to Y_{R(c)}$. Moreover, we can select the $Y_{1,c}$ and $Y_{c}$ so that $Y_{c}\Supset Y_{1,c}$.
@@
[…]
@@ 3. $\psi^{\bullet}$-ql renormalization and near-degenerate regime
Given a compact subset $K\Subset S$, we denote by ${\mathcal{F}}(S,K)$ the family of non-trivial proper curves in $S\setminus K$ connecting $\partial K$ to $\partial S$. We write
@@ 3.1. Outline
The correspondence $F_{n,i}$ in (3.3) is called a $\psi^{\bullet}$-ql map. It consists of a degree $2$ branched covering $f_{n,i}$ and an immersion $\iota_{n,i}$, see definitions in §3.4.2. We define the Julia set and bush of $F_{n,i}$ to be that of the ql map $f_{n,i}$ (3.2):
@@
A $\psi^{\bullet}$-ql renormalization can be naturally defined for a $\psi^{\bullet}$-ql map; i.e., $\psi^{\bullet}$-ql renormalization can be iterated. The Sup-Chain Rule for renormalization domains is stated in §3.4.4.
@@
by considering the fiber product §3.4.6. We have a natural $2^{k}:1$ branched covering $f^{k}\colon U^{k}\to V$ and an immersion $\iota^{k}\colon U^{k}\to V$.
@@ A Euclidean rectangle is a rectangle $E_{x}\coloneqq[0,x]\times[0,1]\subset{\mathbb{C}}$, where:
Assume that ${\mathcal{W}}(E_{x})>2$. The left and right $1$-buffers of $E_{x}$ are defined
@@
is obtained from $E_{x}$ by removing $1$-buffers. If ${\mathcal{W}}(E_{x})\leq 2$, then we set $E_{x}^{\operatorname{new}}\coloneqq\emptyset$. Similarly, buffers of any width are defined.
@@ 3.2.1. Monotonicity and Grötzsch inequality
We say a family of curves $S\displaystyle\SS$ overflows a family ${\mathcal{G}}$ if every curve $S\displaystyle\gamma\in\SS$ contains a subcurve $\gamma^{\prime}\in{\mathcal{G}}$. We also say that
@@ 3.3.1. Proper rectangles
A proper rectangle $R$ in an open hyperbolic surface $S$ is a rectangle in its ideal completion $\overline{S}=S\cup\partial^{i}S$ such that $\partial^{h}R\subset\partial^{i}S$. We naturally view vertical curves of $R$ as proper (and open) paths in $S$. We say that $R$ connects boundary components $J_{0},J_{1}\subset\partial^{i}S$ if $\partial^{h,0}R\subset J_{0},\ \partial^{h,1}R\subset J_{1}$.
@@ 3.3.2. Universal cover
For compact Riemann surface $V$ as in §2.4.1, write its non-empty boundary as $\partial V=\bigsqcup_{k}J_{k}$. We assume that $V$ has negative Euler characteristic $\chi(V)<0$. We set $\pi:\mathbb{D}\rightarrow\operatorname{int}V$ to be the universal covering, and we let ${\Lambda}\subset{\mathbb{T}}=\partial\mathbb{D}$ be the limit set for the group ${\Delta}$ of deck transformations of $\pi$. Then $\pi$ extends continuously to ${\mathbb{T}}\setminus{\Lambda}$, and $\pi:\overline{{\mathbb{D}}}\setminus{\Lambda}\rightarrow V$ is the universal covering of $V$. Moreover, $\pi$ restricted to any component of ${\mathbb{T}}\setminus{\Lambda}$ gives us a universal covering of some component $J_{k}$ of $\partial V$. Such a component of ${\mathbb{T}}\setminus{\Lambda}$ will be denoted ${\mathbf{J}}_{k}$ (usually we will need only one component for each $k$, so we will not use an extra label for it). The stabilizer of ${\mathbf{J}}_{k}$ in ${\Delta}$ is a cyclic group generated by a hyperbolic Möbius transformation $\tau_{k}$ with fixed points in $\partial{\mathbf{J}}_{k}$.